Operation Summary for February 1, 2010 - February 3, 2010 |
Monday Feb. 1 The day shift began with Operations monitoring the Tevatron (TeV) store 7562 (167.32E30) and the stash (115.6E10), with the Antiproton Source stacking (167.37E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations halted stacking at 9:30 AM for a water cage access to investigate a return air pressure issue with the AP0 dump; Pbar resumed stacking at 11 AM. Store 7562 collided until the sector B11 vertical separator sparked at noon causing a quench at E1. TeV personnel investigated the separator problems through the end of the shift. They accessed the TeV A-E areas for an inspection. A Linac Klystron RF station (KRF4) solenoid power supply failed at 11:15 AM. Linac personnel replaced the supply and Operations resumed sending beam at 2:30 PM. During the evening shift, the TeV A-E access concluded at 5:30 PM; operators began the TeV turn on and checkout. Operations put the TeV into HEP shot setup at 10:15 PM. Shot setup ended prematurely with a quench at A1 during step 14 of the squeeze. The quench investigation continued through to the end of shift. Tuesday Feb. 2 The midnight shift began with a continuing TeV quench investigation, with Operations monitoring the stash (65.8E10), with Pbar stacking (8.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators began the TeV and cryo system recovery. The A11 Kautzky valve did not close following the quench. Cryo system technicians accessed the Transfer Hall to reseat the valve. TeV turn-on followed and the machine was ready to operate by 6 AM. A Linac Klystron RF station (KRF1) developed a red heat filament problem at 5 AM that required a time-in period to come back up to temperature. Operations reported that KRF1 was quite unstable following the filament time-in period and required attention from Linac personnel. During the day shift, Linac personnel continued to retune KRF1 following the red heat time-in period. Linac resumed sending beam at 9:30 AM. TeV personnel conducted parsing studies all shift in order to correct beam positions through the squeeze. During the evening shift, Operations put the TeV into shot setup at 4:30 PM. Operators noticed a noisy regulator for a power supply (HF24); EE Support personnel changed it out. Operations established store 7573 at 6:30 PM with an Average Initial Luminosity (AIL) of 243.9E30. The interlock module for a Pbar Accumulator kicker (A:EKIK) failed, causing minor problems and delays with antiproton transfers to the Recycler. Wednesday Feb. 3 The midnight shift began with Operations monitoring store 7573 (109.53E30) and the stash (158.2E10), with Pbar stacking (19.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 5:15 AM, store 7573 aborted when a TeV RF station (TRF2) tripped off due to repeated gap faults. TRF4 had tripped off earlier in the shift. Operators prepared the TeV for D0 and A-E accesses, which began at 6:30 AM. Operations monitored the MI RF vacuum. The Plans for Wednesday The plans for today are to complete the TeV accesses and then resume normal stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 02/03/2010 email Fermilab |
| Security, Privacy, Legal | |